# Guided Policy Optimization under Partial Observability

### v1

@@ Appendix A Related Works
Although leveraging historical information has proven effective for solving POMDPs [14, 25, 21], additional information—often available in simulators during training—can be exploited to further aid learning. Leveraging additional information to accelerate learning in POMDPs has been explored across various frameworks and application domains [47, 16, 17]. A prominent line of research focuses on Imitation Learning (IL), where expert knowledge, often equipped with extra information, significantly enhances performance in practical domains like autonomous driving [3, 9] and robot navigation and planning [7, 4]. However, traditional IL methods such as Behavioral Cloning (BC) [32, 46] and DAgger [34] often lead to sub-optimal solutions in scenarios requiring active information gathering by the agent [31, 49].
@@
To overcome these limitations, recent research has focused on hybrid approaches that integrate RL with IL, often in the context of policy distillation [8]. For instance, [29] modifies Soft Actor Critic (SAC) [12] by replacing the entropy term with a divergence measure between agent and expert policies at each visited state. Similarly, [50] introduces a balancing mechanism between BC and RL training, adjusting based on the agent’s ability to mimic the expert. Additionally, [48] applies potential-based reward shaping [28] using the expert’s value function to guide the agent’s policy gradient, while [40] augments entropy in SAC to blend task reward with expert guidance, where the balance is based on the agent’s performance relative to a reward-only learner.
@@
Despite these advances, expert-driven approaches often assume access to a reliable expert, which may not be feasible when only supplementary information is available. This has led to a growing body of work on co-training approaches where the expert and agent are learned jointly, with the expert conditioned on additional information. For example, [35] proposes training separate policies for the agent and expert using spatial attention for image-based RL, aligning attention mechanisms through shared experiences. [41] co-trains two policies, each conditioned on different information, and selects the most successful rollouts from both policies to guide subsequent learning via RL or IL. [49] further develops this idea in adaptive asymmetric DAgger (A2D), where the expert is continuously refined through RL while supervising the agent. [51] also co-trains a teacher and a student, using the experience collected by teacher to apply RL and BC for the student. Beyond expert-based methods, a complementary approach involves embedding supplementary information directly into the value function within the actor-critic framework [31, 1, 2], which is also called asymmetric learning. This approach is particularly useful in multi-agent settings where global information is naturally accessible [10, 22, 54]. Additional strategies include learning from noisy demonstrations [44], improving via self-correction from past trajectories [42], and surpassing imperfect experts through regret-minimization frameworks [5]. Recent work also explores leveraging LLMs as privileged experts, such as in embodied agents trained with reflective text-based guidance [53]. In our experiments, we benchmark against several algorithms inspired by these lines of work, with detailed descriptions of the baselines provided in Appendix E.1.
@@ Appendix B Omitted Proofs
$\hat{\mu}=\text{arg}\min\{-\eta_{k}\langle\nabla V(\mu^{(k)}),\mu\rangle+\frac{1}{1-\gamma}D_{\mu^{(k)}}(\mu,\mu^{(k)})\},$ (13)
@@
$\pi^{(k+1)}=\underset{\pi\in\Pi}{\text{arg}\min}\{-\eta_{k}\langle\nabla V(\pi^{(k)}),\pi\rangle+\frac{1}{1-\gamma}D_{\pi^{(k)}}(\pi,\pi^{(k)})\}$ (14)
@@
$D_{q}(p,q)=h_{q}(p)-h_{q}(q)-\langle\nabla h_{q}(q),p-q\rangle,$ (15)
@@
$\hat{\mu}$ $=\text{arg}\min\bigg{\{}-\eta_{k}\langle\nabla V(\mu^{(k)}),\mu\rangle+\frac{1}{1-\gamma}D_{\mu^{(k)}}(\mu,\mu^{(k)})\bigg{\}}$ (16)
$=\text{arg}\min\bigg{\{}-\eta_{k}\langle\nabla V(\pi^{(k)}),\mu\rangle+\frac{1}{1-\gamma}D_{\pi^{(k)}}(\mu,\pi^{(k)})\bigg{\}}$
@@
$-(1-\gamma)\eta_{k}\nabla V(\mu^{(k)})+\nabla h_{\mu^{(k)}}(\hat{\mu})-\nabla h_{\mu^{(k)}}(\mu^{(k)})=0,$ (17)
@@
$\nabla_{p}D_{q}(p,q)=\nabla_{p}h_{q}(p)-\nabla_{p}h_{q}(q).$ (18)
@@
$=\underset{\pi\in\Pi}{\text{arg}\min}\{-\eta_{k}\langle\nabla V(\pi^{(k)}),\pi\rangle+\frac{1}{1-\gamma}D_{\pi^{(k)}}(\pi,\pi^{(k)})\}$
@@
For policy $\pi$, $\mu$, $\beta$ and all state $s$, suppose $D_{TV}(\mu(\cdot|s),\beta(\cdot|s))\lesssim\epsilon/2$, then we have
@@
First, let’s examine the assumption $D_{TV}(\mu(\cdot|s),\beta(\cdot|s))\lesssim\epsilon/2$ to check its validity.
@@
$D_{TV}(\mu(\cdot|s),\beta(\cdot|s))=\frac{1}{2}\sum_{a}|\mu(a|s)-\beta(a|s)|=\frac{1}{2}\sum_{a}\beta(a|s)|\rho^{\mu}(s,a)-1|\lesssim\epsilon/2.$ (22)
This confirms that the assumption $D_{TV}(\mu(\cdot|s),\beta(\cdot|s))\lesssim\epsilon/2$ is reasonable, especially for the first few policy updates.
@@
$D_{TV}(\pi(\cdot|o),\beta(\cdot|s))\leq D_{TV}(\pi(\cdot|o),\mu(\cdot|s))+D_{TV}(\mu(\cdot|s),\beta(\cdot|s)),$ (23)
@@
$D_{TV}(\pi(\cdot|o),\beta(\cdot|s))$ $\leq\sqrt{\frac{1}{2}D_{\text{KL}}(\pi(\cdot|o),\mu(\cdot|s))}+D_{TV}(\mu(\cdot|s),\beta(\cdot|s))$
@@
$\mathbb{E}_{a\sim\beta}\big{[}|1-\rho^{\pi}(s,a)|\big{]}=2D_{TV}(\pi(\cdot|o),\beta(\cdot|s))\lesssim\epsilon+\sqrt{2d_{targ}}.$ (24)
@@ Appendix D GPO on TigerDoor-alt Problem
We provide an intuitive example to illustrate how GPO can achieve the optimal policy in the TigerDoor-alt problem. At time step $t$, suppose the guider’s and learner’s policies are:
@@
where the two policies are equal due to the backtracking from time step $t-1$. After one update step, the guider’s policy becomes:
@@ E.1 Baselines
$\mathcal{L}_{\text{GPO-naive}}(\theta)=-\mathbb{E}\Big{[}\min\Big{(}\rho^{\mu_{\theta}}$ $A^{\beta}(o_{g},a),\rho_{clip}^{\mu_{\theta}}A^{\beta}(o_{g},a)\Big{)}-\alpha\text{D}_{\text{KL}}\big{(}\mu_{\theta}(\cdot|o_{l})||\pi_{\hat{\theta}}(\cdot|o_{g})\big{)}$ (26)
$-\text{D}_{\text{KL}}\big{(}\mu_{\hat{\theta}}(\cdot|o_{l})||\pi_{\theta}(\cdot|o_{g})\big{)}\Big{]}.$
@@
$\mathcal{L}_{\text{GPO-ablation}}(\theta)=-\mathbb{E}\Big{[}\min\Big{(}$ $\rho^{\mu_{\theta}}A^{\beta}(o_{g},a),\rho_{clip}^{\mu_{\theta}}A^{\beta}(o_{g},a)\Big{)}-\alpha\text{D}_{\text{KL}}\big{(}\mu_{\theta}(\cdot|o_{l})||\pi_{\hat{\theta}}(\cdot|o_{g})\big{)}$ (27)
@@
ADVISOR. Given teacher’s policy $\mu$, ADVISOR [50] uses a balancing coefficient $w$ between BC and RL training, based on the distance between the teacher’s policy $\mu$ and an auxiliary imitation policy $\hat{\pi}$:
$\mathcal{L}(\pi)=-\mathbb{E}\bigg{[}w\text{CE}(\mu(\cdot|o_{g}),\pi(\cdot|o_{l}))+(1-w)\min\bigg{(}\rho^{\pi}(o_{l},a)A^{\beta}(o_{l},a),\rho_{clip}^{\pi}(o_{l},a,\epsilon)A^{\beta}(o_{l},a)\bigg{)}\bigg{]},$
where $w=exp(-\alpha D_{\text{KL}}(\mu(\cdot|o_{g}),\hat{\pi}(\cdot|o_{l})))$ and CE means cross-entropy.
@@
A2D. Adaptive Asymmetric DAgger (A2D) [49] is closely related to GPO, as it also involves co-training both the teacher and the student. A2D uses a mixture policy $\beta(a|o_{g},o_{l})=\lambda\mu(a|o_{g})+(1-\lambda)\pi(a|o_{l})$ to collect trajectories and train the expert $\mu$ with a mixed value function $V(o_{g},o_{l})=\lambda V^{\mu}(o_{g})+(1-\lambda)v^{\pi}(o_{l})$. The objective is:
@@
ELF. Given teacher’s policy $\mu$, ELF Distillation [48] trains two policies jointly: a follower $\pi_{f}$ to mimic the teacher through BC:
@@
To utilize teacher supervision, ELF applies a potential-based reward shaping [28] $r+\gamma V^{\pi_{f}}(o_{l}^{\prime})-V^{\pi_{f}}(o_{l})$ to the explorer, where $V^{\pi_{f}}$ is the value function of follower. However, ELF needs to divide the interaction equally to train the follower and the explorer, which leads to inefficiencies.
@@ E.2 Hyperparameters
The experiments in Sections 4.1 and 4.3 use the same codebase from [23]. The hyperparameters for these experiments are listed in Table 5. For GPO-clip, due to the asymmetry with large $\delta$, we replace the clip$(\frac{\mu}{\pi},1-\delta,1+\delta)$ with clip$(\frac{\mu}{\pi},\frac{1}{\rho},\rho)$ in the POPGym tasks.
For the experiments in Section 4.2, we use the codebase from [11]. We perform a hyperparameter search for the original versions of the tasks and then fix the same hyperparameters for the partially observable and noisy variants. The hyperparameter search is detailed in Table 6, and the selected hyperparameters for the experiments are provided in Table 7. Other fixed hyperparameters are listed in Table 8.
@@ E.4 Additional Comparative Experiments
Second, we provide a more detailed comparison with methods that train a teacher first and then apply TSL techniques to address the challenge of an "impossibly good" teacher. Two state-of-the-art approaches in this category are ADVISOR and TGRL, which are based on PPO and SAC, respectively. We evaluate these methods using a PPO-trained teacher with full observability on the Ant task. The results, shown in Fig. 9, indicate that while both methods perform well and demonstrate improved efficiency over their respective base algorithms under full observability, their performance degrades in the partially observable Ant task. In this case—where the teacher’s policy is effectively inimitable—the TSL methods perform comparably to their base RL algorithms.
@@
In addition, we provide the results of TGRL on 28 Brax tasks we used in Fig. 11. Note that TGRL utilizes an off-policy training style, differing from all other methods included in the main paper, making it significantly slower (see Table 9), so we only run it for 1M steps, which the curve shows it already converges.
Figure 11: TGRL with pre-trained teacher on 28 Brax tasks.

### v2

@@ 5 Conclusion and Future work
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China [grant numbers U22A2062, U23B2037, 12272008, 62450001, 62476008].
@@ Appendix A Related Works
Although leveraging historical information has proven effective for solving POMDPs (Igl et al., 2018; Meng et al., 2021; Liu et al., 2022), additional information—often available in simulators during training—can be exploited to further aid learning. Leveraging additional information to accelerate learning in POMDPs has been explored across various frameworks and application domains (Vapnik & Vashist, 2009; Lambert et al., 2018; Lee et al., 2023). A prominent line of research focuses on Imitation Learning (IL), where expert knowledge, often equipped with extra information, significantly enhances performance in practical domains like autonomous driving (Bansal et al., 2018; De Haan et al., 2019) and robot navigation and planning (Choudhury et al., 2017; Bhardwaj et al., 2017). However, traditional IL methods such as Behavioral Cloning (BC) (Pomerleau, 1991; Torabi et al., 2018) and DAgger (Ross et al., 2011) often lead to sub-optimal solutions in scenarios requiring active information gathering by the agent (Pinto et al., 2018; Warrington et al., 2020).
@@
To overcome these limitations, recent research has focused on hybrid approaches that integrate RL with IL, often in the context of policy distillation (Czarnecki et al., 2019). For instance, (Nguyen et al., 2022) modifies Soft Actor Critic (SAC) (Haarnoja et al., 2018) by replacing the entropy term with a divergence measure between agent and expert policies at each visited state. Similarly, (Weihs et al., 2024) introduces a balancing mechanism between BC and RL training, adjusting based on the agent’s ability to mimic the expert. Additionally, (Walsman et al., 2023) applies potential-based reward shaping (Ng et al., 1999) using the expert’s value function to guide the agent’s policy gradient, while (Shenfeld et al., 2023b) augments entropy in SAC to blend task reward with expert guidance, where the balance is based on the agent’s performance relative to a reward-only learner.
@@
Despite these advances, expert-driven approaches often assume access to a reliable expert, which may not be feasible when only supplementary information is available. This has led to a growing body of work on co-training approaches where the expert and agent are learned jointly, with the expert conditioned on additional information. For example, (Salter et al., 2021) proposes training separate policies for the agent and expert using spatial attention for image-based RL, aligning attention mechanisms through shared experiences. (Song et al., 2020) co-trains two policies, each conditioned on different information, and selects the most successful rollouts from both policies to guide subsequent learning via RL or IL. (Warrington et al., 2020) further develops this idea in adaptive asymmetric DAgger (A2D), where the expert is continuously refined through RL while supervising the agent. (Wu et al., 2024) also co-trains a teacher and a student, using the experience collected by teacher to apply RL and BC for the student. Beyond expert-based methods, a complementary approach involves embedding supplementary information directly into the value function within the actor-critic framework (Pinto et al., 2018; Andrychowicz et al., 2020; Baisero & Amato, 2021), which is also called asymmetric learning. This approach is particularly useful in multi-agent settings where global information is naturally accessible (Foerster et al., 2018; Lowe et al., 2017; Yu et al., 2022). Additional strategies include learning from noisy demonstrations (Tangkaratt et al., 2021), improving via self-correction from past trajectories (Song et al., 2018), and surpassing imperfect experts through regret-minimization frameworks (Chang et al., 2015). Recent work also explores leveraging LLMs as privileged experts, such as in embodied agents trained with reflective text-based guidance (Yang et al., 2024).
Besides, there are also representation learning techniques provided in order to reconstruct the privileged information (or its latent representation) via partial observation. For example Sermanet et al. (2018); Seo et al. (2023) use multi-view setups (e.g., image-based manipulation with additional camera views) to learn more informative embeddings. Others (Lee et al., 2020b; Salter et al., 2021; Kumar et al., 2021; Qi et al., 2023) leverage privileged simulator states during training and design policies that operate on both observed and inferred states.
In our experiments, we benchmark against several algorithms inspired by these lines of work, with detailed descriptions of the baselines provided in Appendix E.1.
@@ Appendix B Omitted Proofs
$\hat{\mu}=\text{arg}\min\{-\eta_{k}\langle\nabla V(\mu^{(k)}),\mu\rangle+\frac{1}{1-\gamma}\text{D}_{\mu^{(k)}}(\mu,\mu^{(k)})\},$ (13)
@@
$\pi^{(k+1)}=\underset{\pi\in\Pi}{\text{arg}\min}\{-\eta_{k}\langle\nabla V(\pi^{(k)}),\pi\rangle+\frac{1}{1-\gamma}\text{D}_{\pi^{(k)}}(\pi,\pi^{(k)})\}$ (14)
@@
$\text{D}_{q}(p,q)=h_{q}(p)-h_{q}(q)-\langle\nabla h_{q}(q),p-q\rangle,$ (15)
@@
$\hat{\mu}$ $=\text{arg}\min\bigg\{-\eta_{k}\langle\nabla V(\mu^{(k)}),\mu\rangle+\frac{1}{1-\gamma}\text{D}_{\mu^{(k)}}(\mu,\mu^{(k)})\bigg\}$ (16)
$=\text{arg}\min\bigg\{-\eta_{k}\langle\nabla V(\pi^{(k)}),\mu\rangle+\frac{1}{1-\gamma}\text{D}_{\pi^{(k)}}(\mu,\pi^{(k)})\bigg\}$
@@
$(1-\gamma)\eta_{k}\nabla V(\mu^{(k)})+\nabla h_{\mu^{(k)}}(\hat{\mu})-\nabla h_{\mu^{(k)}}(\mu^{(k)})=0,$ (17)
@@
$\nabla_{p}\text{D}_{q}(p,q)=\nabla_{p}h_{q}(p)-\nabla_{p}h_{q}(q).$ (18)
@@
$=\underset{\pi\in\Pi}{\text{arg}\min}\{-\eta_{k}\langle\nabla V(\pi^{(k)}),\pi\rangle+\frac{1}{1-\gamma}\text{D}_{\pi^{(k)}}(\pi,\pi^{(k)})\}$
@@
For policy $\pi$, $\mu$, $\beta$ and all state $s$, suppose $\text{D}_{\text{TV}}(\mu(\cdot|s),\beta(\cdot|s))\lesssim\epsilon/2$, then we have
@@
First, let’s examine the assumption $\text{D}_{\text{TV}}(\mu(\cdot|s),\beta(\cdot|s))\lesssim\epsilon/2$ to check its validity.
@@
$\text{D}_{\text{TV}}(\mu(\cdot|s),\beta(\cdot|s))=\frac{1}{2}\sum_{a}|\mu(a|s)-\beta(a|s)|=\frac{1}{2}\sum_{a}\beta(a|s)|\rho^{\mu}(s,a)-1|\lesssim\epsilon/2.$ (22)
This confirms that the assumption $\text{D}_{\text{TV}}(\mu(\cdot|s),\beta(\cdot|s))\lesssim\epsilon/2$ is reasonable, especially for the first few policy updates.
@@
$\text{D}_{\text{TV}}(\pi(\cdot|o),\beta(\cdot|s))\leq D_{TV}(\pi(\cdot|o),\mu(\cdot|s))+\text{D}_{\text{TV}}(\mu(\cdot|s),\beta(\cdot|s)),$ (23)
@@
$\text{D}_{\text{TV}}(\pi(\cdot|o),\beta(\cdot|s))$ $\leq\sqrt{\frac{1}{2}\text{D}_{\text{KL}}(\pi(\cdot|o),\mu(\cdot|s))}+\text{D}_{\text{TV}}(\mu(\cdot|s),\beta(\cdot|s))$
@@
$\mathbb{E}_{a\sim\beta}\big[|1-\rho^{\pi}(s,a)|\big]=2\text{D}_{\text{TV}}(\pi(\cdot|o),\beta(\cdot|s))\lesssim\epsilon+\sqrt{2d_{targ}}.$ (24)
@@ Appendix D GPO on TigerDoor-alt Problem
Here we provide an intuitive example to illustrate how GPO can achieve the optimal policy in the TigerDoor-alt problem. At time step $t$, suppose the guider’s and learner’s policies are:
@@
where the two policies are equal due to the backtracking from time step t1t-1. After one update step, the guider’s policy becomes:
@@ E.1 Baselines
$\mathcal{L}_{\text{GPO-naive}}(\theta)=-\mathbb{E}\Big[\min\Big(\rho^{\mu_{\theta}}$ $A^{\beta}(o_{g},a),\rho_{clip}^{\mu_{\theta}}A^{\beta}(o_{g},a)\Big)-\alpha\text{D}_{\text{KL}}\big(\mu_{\theta}(\cdot|o_{l}),\pi_{\hat{\theta}}(\cdot|o_{g})\big)$ (26)
$-\text{D}_{\text{KL}}\big(\mu_{\hat{\theta}}(\cdot|o_{l}),\pi_{\theta}(\cdot|o_{g})\big)\Big].$
@@
$\mathcal{L}_{\text{GPO-ablation}}(\theta)=-\mathbb{E}\Big[\min\Big($ $\rho^{\mu_{\theta}}A^{\beta}(o_{g},a),\rho_{clip}^{\mu_{\theta}}A^{\beta}(o_{g},a)\Big)-\alpha\text{D}_{\text{KL}}\big(\mu_{\theta}(\cdot|o_{l}),\pi_{\hat{\theta}}(\cdot|o_{g})\big)$ (27)
@@
ADVISOR. Given teacher’s policy $\mu$, ADVISOR (Weihs et al., 2024) uses a balancing coefficient $w$ between BC and RL training, based on the distance between the teacher’s policy $\mu$ and an auxiliary imitation policy $\hat{\pi}$:
$\mathcal{L}(\pi)=-\mathbb{E}\bigg[w\text{CE}(\mu(\cdot|o_{g}),\pi(\cdot|o_{l}))+(1-w)\min\Big(\rho^{\pi}(o_{l},a)A^{\beta}(o_{l},a),\rho_{clip}^{\pi}(o_{l},a,\epsilon)A^{\beta}(o_{l},a)\Big)\bigg],$
where $w=exp(-\alpha\text{D}_{\text{KL}}(\mu(\cdot|o_{g}),\hat{\pi}(\cdot|o_{l})))$ and CE means cross-entropy.
@@
A2D. Adaptive Asymmetric DAgger (A2D) (Warrington et al., 2020) is closely related to GPO, as it also involves co-training both the teacher and the student. A2D uses a mixture policy $\beta(a|o_{g},o_{l})=\lambda\mu(a|o_{g})+(1-\lambda)\pi(a|o_{l})$ to collect trajectories and train the expert $\mu$ with a mixed value function $V(o_{g},o_{l})=\lambda V^{\mu}(o_{g})+(1-\lambda)v^{\pi}(o_{l})$. The objective is:
@@
ELF. Given teacher’s policy $\mu$, ELF Distillation (Walsman et al., 2023) trains two policies jointly: a follower $\pi_{f}$ to mimic the teacher through BC:
@@
To utilize teacher supervision, ELF applies a potential-based reward shaping (Ng et al., 1999) $r+\gamma V^{\pi_{f}}(o_{l}^{\prime})-V^{\pi_{f}}(o_{l})$ to the explorer, where $V^{\pi_{f}}$ is the value function of follower. However, ELF needs to divide the interaction equally to train the follower and the explorer, which leads to inefficiencies.
@@ E.2 Hyperparameters
[…]
For the experiments in Section 4.2, we use the codebase from (Freeman et al., 2021). We perform a hyperparameter search for the original versions of the tasks and then fix the same hyperparameters for the partially observable and noisy variants. The hyperparameter search is detailed in Table 6, and the selected hyperparameters for the experiments are provided in Table 7. Other fixed hyperparameters are listed in Table 8.
All algorithms in Brax are run with 10 random seeds, whereas those in POPGym use 3 seeds, as the latter exhibits lower variance. Reward curves in this paper report the mean and standard deviation across these runs.
@@ E.4 Additional Comparative Experiments
Second, we provide a more detailed comparison with methods that train a teacher first and then apply TSL techniques to address the challenge of an inimitable teacher. Two state-of-the-art approaches in this category are ADVISOR and TGRL, which are based on PPO and SAC, respectively. We evaluate these methods using a PPO-trained teacher with full observability on the Ant task. The results, shown in Fig. 9, indicate that while both methods perform well and demonstrate improved efficiency over their respective base algorithms under full observability, their performance degrades in the partially observable Ant task. In this case—where the teacher’s policy is effectively inimitable—the TSL methods perform comparably to their base RL algorithms.
@@
Finally, we include the representation learning method RMA (Kumar et al., 2021), which aims to reconstruct the latent privileged information used by the teacher during the student training phase. Such representation learning approaches are promising when privileged information can be reliably approximated from partial observations. However, their effectiveness is limited in Brax, where observations are noisy (Figure 11). Since the noise cannot be removed by a deterministic mapping without additional structure or assumptions, regression-based reconstruction tends to collapse to identity mappings and fails to recover meaningful latent representations. Moreover, current encoder-based pipelines often lack theoretical guarantees for convergence or generalization across diverse tasks, particularly under partial observability.
@@
Figure 11: TGRL and RMA with pre-trained teacher on 28 Brax tasks.